OpenStudy (anonymous):

rewrite sin^3x-cos^2x sin^3x as a single trigonometric function

OpenStudy (gorv):

\[\sin^3x-\cos^2x *\sin^3x=\sin^3x(1-\cos^2x)=\sin^3x * \sin^2x= \sin^5x\]

OpenStudy (gorv):

take sin^3x common and 1-cos^2x=sin^2x

OpenStudy (gorv):

@kierra21